Shift report 20030512 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon May 12 2003 - 15:27:45 EDT

  • Next message: Flemming Videbaek: "near term scheduling"
    ------------------------------------------------------------
    Date: Mon May 12 15:27:45 2003, Shift: 08:00-16:00
    Supervisor: fv, with: 
    Log book#14 pages 108 - 108
    
    Shift Summary:
    ------------------------------------------------------------
    1) 
    Beam development restoring polarization which succeeded. 
    Beam available from ~ 12.15 albeit only in physics mode much later. 
    The big boys voted for dumping - dumped at 15:15 
     
    2)  
    Tdc calibration run 8919 
    Decided to call 1/2A 12 deg for complete 3.8 K trigger 2. 
    After the 4 deg high filed and pol running 12 deg can be cobntinued at 2/3 
    field to optimize pt ~3 GeV/c. 
     
    Data run 8925,8926,8927  100K trigger 6 12 deg at 1/15 field. 
    zero filed run 8928  (2 k trigger 6) needs more. 
    Worked to implement more (last) input to spin scalers with Brendan Fox. 
    (ZDC gates narrow+ Background Blue,yellow) 
     
    3) No new problem  - attempted to dix some bad tofw and tfw2 slats 
      see e-log for details. 
     
    A: we now attempt to start and stop the spin scalers on a routine basis 
    i.e. RSstart.perl before run and RSstop.perl after run. 
     
     
    Store hours 
    took data from 12:30 to 15:15 == 2.75 hours 
    
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Mon May 12 2003 - 15:28:36 EDT